a group g is isomorphism to one of its proper subgroup then g=z is it true z=integers geoup


1
Expert's answer
2021-04-15T07:27:54-0400

No, it is false.

As a counter-example we can think of a group "\mathbb{Z}^2". It is isomorphic to the group "2\mathbb{Z} \times \mathbb{Z}", as "\mathbb{Z} \simeq 2\mathbb{Z}" which is a proper subgroup. At the same time "\mathbb{Z}^2 \nsim \mathbb{Z}", as "\mathbb{Z}" has one generator and "\mathbb{Z}^2" has two of them.